Two cars leave town going in the same direction on the same road. One travels at 30mph and the other at 46mph. In how many hours will they be 72 miles apart?
Answer by checkley75(3666)

46X-30X=72
16X=72
X=72/16
X=4.5 HOURS & THEY WILL BE 72 MILES APART.
PROOF
46*4.5-30*4.5=72
207-135= 72
72=72